# RealEngine: Simulating Autonomous Driving in Realistic Context

### v1

@@ 3.1.3 Driving scenario composition
The combination of reconstructed background scenes and traffic participant models forms a rich design space for creating diverse and flexible driving scenarios. Scenario composition begins by selecting a background scene, into which a set of traffic participants is inserted. Each participant is assigned an initial position, orientation, and free-form trajectory, either manually specified or dynamically planned by a driving agent. These participants are then spatially registered into the background scene’s coordinate system. An overview is shown in Figure 2.
To ensure realistic integration between foreground participants and background scenes, we introduce a physically based rendering (PBR) process that optimizes environment light maps for consistent relighting and shadow casting. We adopt the Disney BRDF lighting model wang2023fegr , where the foreground color is computed as:
@@
where ${\bm{C}}_{bg}$ is the background image, ${\bm{C}}_{fg}$ and ${\bm{M}}_{fg}$ are the rendered foreground RGB and mask, respectively. We optimize $\{{\bm{L}}_{i},{\bm{L}}_{s}\}$ using StableDiffusion sd with SDS loss poole2022dreamfusion , ensuring photorealistic blending.
@@ 3.2 Closed-loop driving simulation
Built upon this flexible scenario composition and novel view synthesis-enabled reconstruction, RealEngine supports a wide range of driving scenarios. We demonstrate three representative categories:
Non-reactive Simulation. Evaluates closed-loop planning where other participants follow fixed pre-recorded trajectories (Figure 1 (a)). The agent perceives rendered multi-modal sensor data, plans trajectories, and executes motion via a linear quadratic regulator (LQR), completing the closed-loop cycle.
Safety Test Simulation. Evaluates agent reactions to inserted participants exhibiting hazardous behaviors (e.g., sudden lane changes or blocked intersections), testing safety-critical skills (Figure 1 (b)).
Multi-agent Interaction Simulation. Simulates cooperative and adversarial interactions where multiple agents, each running independent planning, simultaneously navigate the scene (Figure 1 (c)).
@@ 3.3 Assessment of driving agents
Driving agent assessment aims to evaluate the agent’s trajectory planning performance across both space and time, considering safety, progress, and interaction quality within diverse traffic scenarios. In traditional open-loop evaluation, the Predictive Driver Model Score (PDMS) Dauner2024NEURIPS measures trajectory quality at individual time steps. However, open-loop PDMS does not capture how planning decisions evolve over time, nor does it account for interactions with dynamic participants in the scene.
@@
PVG chen2023periodic OmniRe chen2024omnire
@@
StreetGaussians yan2024street RealEngine (Ours)
@@
Figure 3: (a) Compared to state-of-the-art reconstruction methods, we achieve superior camera image reconstruction in the nuPlan nuplan benchmark. Additionally, (b) our foreground relighting technique enables seamless integration of inserted participants with the background scene.
(a) ST-P3 hu2022stp3
(b) DiffusionDrive liao2024diffusiondrive
Figure 4: Non-reactive simulation. The driving agent’s planned trajectory at each step is visualized in bird’s-eye view (blue) and the front-view camera (green). During closed-loop simulation, ST-P3 hu2022stp3 exhibits inconsistencies in consecutive frame planning, leading to error accumulation and causing the vehicle to navigate into invalid regions. In contrast, DiffusionDrive liao2024diffusiondrive maintains more consistent planning across consecutive frames, resulting in a higher PDM Score.
@@ 4 Experiments
Datasets. We use CO3D reizenstein21co3d and 3DRealCar du20243drealcar to reconstruct diverse foreground traffic participants, followed by mesh extraction for rendering and scene composition. Additionally, we include high-quality meshes from Sketchfab Sketchfab for further diversity in foreground insertion and relighting.
For background scene reconstruction and scene editing, we leverage Navsim Dauner2024NEURIPS , which is derived from OpenScene openscene2023 , a simplified version of nuPlan nuplan . We select 14 diverse sequences from Navsim for scene editing and driving agent evaluation, and design 14, 21, and 14 test cases for the non-reactive, safety, and multi-agent interaction simulation, respectively. To ensure high-quality sensor simulation, we retrieve high-frequency images and LiDAR point clouds from the corresponding nuPlan sequences for background scene reconstruction. Using an NVIDIA RTX A6000, the rendering frame rate reaches 30Hz for camera images and 15Hz for LiDAR data.
Autonomous driving models. We evaluate four end-to-end driving models: ST-P3 hu2022stp3 , VAD jiang2023vad , TransFuser Chitta2023PAMI , and DiffusionDrive liao2024diffusiondrive , with ST-P3 and VAD reimplemented on nuPlan nuplan for consistency. As a baseline, we also test Navsim’s constant velocity model for comparison.
@@
Constant velocity Dauner2024NEURIPS 92.9 64.3 85.7 100 29.4 46.8
ST-P3 hu2022stp3 Open-loop 92.9 71.4 92.9 100 46.2 59.6
VAD jiang2023vad 92.9 85.7 92.9 100 48.5 66.1
TransFuser Chitta2023PAMI 92.9 85.7 92.9 100 55.9 69.1
DiffusionDrive liao2024diffusiondrive 92.9 85.7 92.9 100 56.7 69.5
ST-P3 hu2022stp3 Closed-loop 100 64.3 85.7 100 35.6 47.5
VAD jiang2023vad 85.7 78.6 78.6 100 34.3 53.0
TransFuser Chitta2023PAMI 92.9 71.4 85.7 100 46.0 57.9
DiffusionDrive liao2024diffusiondrive 92.9 71.4 92.9 100 47.1 61.3
@@
(a) Stationary          (ST-P3 hu2022stp3 ) (b) Sudden brake           (VAD jiang2023vad ) (c) Intersection encounter (TransFuser Chitta2023PAMI ) (d) Cut in (DiffusionDrive liao2024diffusiondrive )
@@ 4.1 Diving scenario quality
We compared our optimized reconstruction results (Appendix A) with the state-of-the-art methods yan2024street ; chen2024omnire ; chen2023periodic on 6 Navsim Dauner2024NEURIPS sequences, as shown in Figure 3 (a) and Table 3, and observed a notable improvement in the reconstruction performance on nuPlan. Meanwhile, our foreground insertion blends more harmoniously with the background and can be recognized by perception models borse2021inverseform ; xie2021segformer ; ren2024grounded , as shown in Figure 3 (b). This enables the driving agents to correctly recognize the inserted foreground participants and respond accordingly.
Table 3: Comparison with state-of-the-art camera images reconstruction methods in the nuPlan benchmark.
[…]
StreetGaussians yan2024street 26.02 0.798 0.179
Omnire chen2024omnire 26.89 0.837 0.165
PVG chen2023periodic 28.32 0.854 0.176
RealEngine(Ours) 29.67 0.897 0.093
(a) Agent 1 (DiffusionDrive liao2024diffusiondrive )
(b) Agent 2 (DiffusionDrive liao2024diffusiondrive )
Figure 6: Multi-agent interaction simulation. The trajectory annotations are consistent with Figure 5. We set two agents (DiffusionDrive liao2024diffusiondrive ) to plan trajectories simultaneously and independently in the showed turning scenario. The first agent avoids collision by increasing its turning radius, while the second agent decelerates to allow safe passage. Once the vehicles have passed each other, the second agent resumes its original speed.
Table 4: Safety test and multi-agent simulation. The notations are consistent with the non-reactive simulation Table 2 above.
@@
Constant velocity Dauner2024NEURIPS Safety test 47.6 71.4 38.1 100 36.7 36.3
ST-P3 hu2022stp3 47.6 100 42.9 100 44.7 44.4
VAD jiang2023vad 47.6 95.2 28.6 100 41.2 37.0
TransFuser Chitta2023PAMI 47.6 100 38.1 100 44.1 42.2
DiffusionDrive liao2024diffusiondrive 57.1 100 52.4 100 54.0 53.8
Constant velocity Dauner2024NEURIPS Multi-agent 42.8 60.7 39.3 100 27.8 27.4
ST-P3 hu2022stp3 53.6 96.4 50.0 100 44.6 46.3
VAD jiang2023vad 32.1 71.4 32.1 100 27.7 28.8
TransFuser Chitta2023PAMI 60.7 96.4 53.6 100 54.3 55.0
DiffusionDrive liao2024diffusiondrive 57.1 96.4 50.0 100 51.7 51.9
@@ 4.2 Closed-loop simulation
Non-reactive simulation. In the non-reactive setting, we conduct open-loop and closed-loop simulations on the driving agents and compare their performance, as shown in Table 2. Compared to directly predicting multiple frame trajectories in open-loop simulation, the PDM Score of the continuously predicted trajectories by the driving agent in closed-loop simulation is reduced to varying degrees. In particular, models tend to exhibit inconsistency in the adjacent predicted trajectories. This inconsistency leads to cumulative errors, causing the vehicle to navigate to unreasonable areas and resulting in a decrease in the DAC (drivable area compliance) metric, as shown in Figure 4.
Safety test simulation. For safety test simulation, we designed three test cases (simple, moderate, and challenging) tailored to each specific scenario and conducted closed-loop simulations to compute the PDM Score. As shown in Figure 5, in simple and moderate scenarios, such as a stationary vehicle blocking the lane or a sudden brake, driving agents can reasonably maneuver to avoid collisions. However, in more challenging scenarios, such as aggressive lane changes or intersection encounters, driving agents may exhibit some reaction but fail to avoid collisions, or in some cases, show minimal response to an imminent collision, leading to a significant drop in the PDM Score, as shown in Table 4.
Multi-agent interaction simulation. In the multi-agent interaction simulation, each vehicle is assigned to an instance of a certain driving agent model for independent and simultaneous control, with only the initial speed and high-level driving commands (right, left, straight, or unknown) specified. For each specific scenario, we design both simple and challenging test cases. At each time step, we render the sensor data for all driving agents to plan trajectories and conduct closed-loop simulations to compute the PDM scores for all instances and average the results, as presented in Table 4. Additionally, Figure 6 illustrates a turning scenario involving two DiffusionDrive liao2024diffusiondrive instances, where both agents successfully avoid collisions by adopting different yet reasonable strategies.
@@ 5 Conclusion
Appendix A Urban scene reconstruction
A.1 NuPlan pose correction
We propose a pose correction method based on LiDAR point cloud registration in Section 3.1.1. As shown in Figure 7, point cloud registration effectively corrects pose calibration errors. This leads to improved scene reconstruction quality, which is reflected in the increased PSNR, as presented in Table 5.
@@ A.1 NuPlan pose correction
A.2 Camera images reconstruction
@@
We use StreetGaussians yan2024street to reconstruct the scene and perform additional processing on the nuPlan dataset. StreetGaussians yan2024street uses manual pose annotation of dynamic objects to distinguish between static background and moving objects. Dynamic objects are reconstructed in their respective centered canonical spaces and then placed into the background scene space during the rendering process based on the known poses.
As shown in Figure 8 (a), the camera images in nuPlan exhibit significant barrel distortion. Since 3DGS relies on image-based rendering and does not render rays corresponding to each pixel as NeRF does, such distortion adversely affects the reconstruction of scene geometry. To mitigate this issue, we applied a distortion correction to the images. Furthermore, due to varying exposure levels across the cameras in the nuPlan dataset, the same object exhibits substantial color differences when viewed from different cameras, as shown in Figure 8 (b), leading to ambiguities in the color information of Gaussian primitives. To address this, we introduced a learnable affine transformation $\{{\bm{A}}_{i},{\bm{t}}_{i}\}$ for each camera $i$ to individually calibrate the image colors:
@@
where $i$ represents the camera index, ${\bm{C}}_{i}$ denotes the original rendered RGB map, and $\tilde{{\bm{C}}}_{i}$ refers to the RGB map after exposure transformation. These processing methods improve the quality of scene reconstruction, as shown in the ablation study in Table 5.
@@
We compared our optimized reconstruction results with the state-of-the-art methods yan2024street ; chen2024omnire ; chen2023periodic on 6 Navsim Dauner2024NEURIPS sequences, as shown in Figure 3 and Table 3, and observed a notable improvement in the reconstruction performance on nuPlan.
@@
Although existing urban reconstruction methods perform excellently in synthesizing novel viewpoints for recorded trajectories, they face challenges when handling new trajectories in closed-loop simulations due to the limited viewpoints of driving videos and the vastness of the driving environment. To address this challenge, we employ DriveX yang2024drivex to enhance the reconstructed scene, which utilizes video generative prior yu2024viewcrafter to optimize the scene model across multiple trajectories. As shown in Figure 9, our final scene model is capable of generating high-fidelity virtual driving environments beyond the recorded trajectories, enabling free-form trajectory driving simulations.
@@
Figure 9: Lane change reconstruction. We render the camera images by shifting the driving perspective 3 meters to the left and right. Due to the limited viewpoint of driving videos and the vastness of the driving environment, large translations from driving perspectives lead to a significant decline in reconstruction quality. To address this, we use DriveX yang2024drivex , which leverages video generative prior to optimize the scene Gaussian primitives, resulting in improved reconstruction quality even with large viewpoint translations.
A.3 LiDAR reconstruction

### v2

@@ 3.1.3 Driving scenario composition
The combination of reconstructed background scenes and traffic participant models forms a rich design space for creating diverse and flexible driving scenarios. Scenario composition begins by selecting a background scene, into which a set of traffic participants is inserted. Each participant is assigned an initial position, orientation, and free-form trajectory, either manually specified or dynamically planned by a driving agent. These participants are then spatially registered into the background scene’s coordinate system. An overview is shown in Figure˜2.
To ensure realistic integration between foreground participants and background scenes, we introduce a physically based rendering (PBR) process that optimizes environment light maps for consistent relighting and shadow casting. We adopt the Disney BRDF lighting model (Wang et al., 2023b), where the foreground color is computed as:
@@
where ${\bm{C}}_{bg}$ is the background image, ${\bm{C}}_{fg}$ and ${\bm{M}}_{fg}$ are the rendered foreground RGB and mask, respectively. We optimize $\{{\bm{L}}_{i},{\bm{L}}_{s}\}$ using StableDiffusion (Rombach et al., 2022) with SDS loss (Poole et al., 2022), ensuring photorealistic blending.
@@ 3.2 Closed-loop driving simulation
Built upon this flexible scenario composition and novel view synthesis-enabled reconstruction, RealEngine supports a wide range of driving scenarios. We demonstrate three categories:
Non-reactive Simulation. Evaluates closed-loop planning where other participants follow fixed pre-recorded trajectories (Figure˜1 (a)). The agent perceives rendered multi-modal sensor data, plans trajectories, and executes motion via a linear quadratic regulator, completing the closed-loop cycle.
Safety Test Simulation. Evaluates agent reactions to inserted participants exhibiting hazardous behaviors (e.g., sudden lane changes or blocked intersections), testing safety-critical skills (Figure˜1 (b)).
Multi-agent Interaction Simulation. Simulates cooperative and adversarial interactions where multiple agents, each running independent planning, simultaneously navigate the scene (Figure˜1 (c)).
@@ 3.3 Assessment of driving agents
Driving agent assessment aims to evaluate the agent’s trajectory planning performance across both space and time, considering safety, progress, and interaction quality within diverse traffic scenarios. In traditional open-loop evaluation, the Predictive Driver Model Score (PDMS) (Dauner et al., 2024) measures trajectory quality at individual time steps. However, open-loop PDMS does not capture how planning decisions evolve over time, nor does it account for interactions with dynamic participants in the scene.
@@
PVG OmniRe
@@
StreetGaussians RealEngine (Ours)
@@
Figure 3: (a) Compared to state-of-the-art reconstruction methods, we achieve superior camera image reconstruction in the nuPlan (H. Caesar, 2021) benchmark. Additionally, (b) our foreground relighting technique enables seamless integration of inserted participants with the background scene.
(a) ST-P3 (Hu et al., 2022)
(b) DiffusionDrive (Liao et al., 2025)
Figure 4: Non-reactive simulation. The driving agent’s planned trajectory at each step is visualized in bird’s-eye view (blue) and the front-view camera (green). During closed-loop simulation, ST-P3 (Hu et al., 2022) exhibits inconsistencies in consecutive frame planning, leading to error accumulation and causing the vehicle to navigate into invalid regions. In contrast, DiffusionDrive (Liao et al., 2025) maintains more consistent planning across consecutive frames, resulting in a higher PDM Score.
@@ 4 Experiments
Datasets. We use CO3D (Reizenstein et al., 2021) and 3DRealCar (Du et al., 2024) to reconstruct diverse foreground traffic participants, followed by mesh extraction for rendering and scene composition. Additionally, we include high-quality meshes from Sketchfab (Contributors, ) for further diversity in foreground insertion and relighting.
For background scene reconstruction and scene editing, we leverage Navsim (Dauner et al., 2024), which is derived from OpenScene (Contributors, 2023), a simplified version of nuPlan (H. Caesar, 2021). We select 14 diverse sequences from Navsim for scene editing and driving agent evaluation, and design 14, 21, and 14 test cases for the non-reactive, safety, and multi-agent interaction simulation, respectively. To ensure high-quality sensor simulation, we retrieve high-frequency images and LiDAR point clouds from the corresponding nuPlan sequences for background scene reconstruction. Using an NVIDIA RTX A6000, the rendering frame rate reaches 30Hz for camera images and 15Hz for LiDAR data.
Autonomous driving models. We evaluate four end-to-end driving models: ST-P3 (Hu et al., 2022), VAD (Jiang et al., 2023), TransFuser (Chitta et al., 2023), and DiffusionDrive (Liao et al., 2025), with ST-P3 and VAD reimplemented on nuPlan (H. Caesar, 2021) for consistency. As a baseline, we also test Navsim’s constant velocity model for comparison.
@@
Constant velocity (Dauner et al., 2024) 92.9 64.3 85.7 100 29.4 46.8
ST-P3 (Hu et al., 2022) Open-loop 92.9 71.4 92.9 100 46.2 59.6
VAD (Jiang et al., 2023) 92.9 85.7 92.9 100 48.5 66.1
TransFuser (Chitta et al., 2023) 92.9 85.7 92.9 100 55.9 69.1
DiffusionDrive (Liao et al., 2025) 92.9 85.7 92.9 100 56.7 69.5
ST-P3 (Hu et al., 2022) Closed-loop 100 64.3 85.7 100 35.6 47.5
VAD (Jiang et al., 2023) 85.7 78.6 78.6 100 34.3 53.0
TransFuser (Chitta et al., 2023) 92.9 71.4 85.7 100 46.0 57.9
DiffusionDrive (Liao et al., 2025) 92.9 71.4 92.9 100 47.1 61.3
@@
(a) Stationary (ST-P3) (b) Sudden brake (VAD) (c) Intersection encounter (TransFuser) (d) Cut in (DiffusionDrive)
@@ 4.1 Diving scenario quality
We compared our optimized reconstruction results (Section˜A.1) with the state-of-the-art methods (Yan et al., 2024b; Chen et al., 2024; 2023) on 6 Navsim (Dauner et al., 2024) sequences, as shown in Figure˜3 (a), and observed a notable improvement in the reconstruction performance on nuPlan. Meanwhile, our foreground insertion blends more harmoniously with the background and can be recognized by perception models (Borse et al., 2021; Xie et al., 2021; Ren et al., 2024), as shown in Figure˜3 (b). This enables the driving agents to correctly recognize the inserted foreground participants and respond accordingly.
(a) Agent 1 (DiffusionDrive (Liao et al., 2025))
(b) Agent 2 (DiffusionDrive (Liao et al., 2025))
Figure 6: Multi-agent interaction simulation. The trajectory annotations are consistent with Figure˜5. We set two agents (DiffusionDrive (Liao et al., 2025)) to plan trajectories simultaneously and independently in the showed turning scenario. The first agent avoids collision by increasing its turning radius, while the second agent decelerates to allow safe passage. Once the vehicles have passed each other, the second agent resumes its original speed.
Table 3: Safety test and multi-agent simulation. The notations are consistent with the non-reactive simulation Table˜2 above.
@@
Constant velocity (Dauner et al., 2024) Safety test 47.6 71.4 38.1 100 36.7 36.3
ST-P3 (Hu et al., 2022) 47.6 100 42.9 100 44.7 44.4
VAD (Jiang et al., 2023) 47.6 95.2 28.6 100 41.2 37.0
TransFuser (Chitta et al., 2023) 47.6 100 38.1 100 44.1 42.2
DiffusionDrive (Liao et al., 2025) 57.1 100 52.4 100 54.0 53.8
Constant velocity (Dauner et al., 2024) Multi-agent 42.8 60.7 39.3 100 27.8 27.4
ST-P3 (Hu et al., 2022) 53.6 96.4 50.0 100 44.6 46.3
VAD (Jiang et al., 2023) 32.1 71.4 32.1 100 27.7 28.8
TransFuser (Chitta et al., 2023) 60.7 96.4 53.6 100 54.3 55.0
DiffusionDrive (Liao et al., 2025) 57.1 96.4 50.0 100 51.7 51.9
@@ 4.2 Closed-loop simulation
Non-reactive simulation. In the non-reactive setting, we conduct open-loop and closed-loop simulations on the driving agents and compare their performance, as shown in Table˜2. Compared to directly predicting multiple frame trajectories in open-loop simulation, the PDM Score of the continuously predicted trajectories by the driving agent in closed-loop simulation is reduced to varying degrees. In particular, models tend to exhibit inconsistency in the adjacent predicted trajectories. This inconsistency leads to cumulative errors, causing the vehicle to navigate to unreasonable areas and resulting in a decrease in the DAC (drivable area compliance) metric, as shown in Figure˜4.
Safety test simulation. For safety test simulation, we designed three test cases (simple, moderate, and challenging) tailored to each specific scenario and conducted closed-loop simulations to compute the PDM Score. As shown in Figure˜5, in simple and moderate scenarios, such as a stationary vehicle blocking the lane or a sudden brake, driving agents can reasonably maneuver to avoid collisions. However, in more challenging scenarios, such as aggressive lane changes or intersection encounters, driving agents may exhibit some reaction but fail to avoid collisions, or in some cases, show minimal response to an imminent collision, leading to a significant drop in the PDM Score, as in Table˜3.
Multi-agent interaction simulation. In the multi-agent interaction simulation, each vehicle is assigned to an instance of a certain driving agent model for independent and simultaneous control, with only the initial speed and high-level driving commands (right, left, straight, or unknown) specified. For each specific scenario, we design both simple and challenging test cases. At each time step, we render the sensor data for all driving agents to plan trajectories and conduct closed-loop simulations to compute the PDM scores for all instances and average the results, as presented in Table˜3. Additionally, Figure˜6 illustrates a turning scenario involving two DiffusionDrive (Liao et al., 2025) instances, where both agents successfully avoid collisions by adopting different yet reasonable strategies.
@@ 5 Conclusion
6 Ethics statement
This work focuses on reconstructing urban scenes and evaluating the performance of autonomous driving models. Our study does not involve human subjects, personally identifiable information, or sensitive private data. All datasets employed are publicly available and widely used in the research community, and we follow the corresponding licenses and usage guidelines. The reconstructed scenes and evaluation results are intended solely for academic research and do not raise immediate safety risks, as the experiments are performed in a simulation environment without deployment in real-world driving systems.
7 Reproducibility statement
We are committed to ensuring the reproducibility of our work. To this end, we will release the full implementation and relevant scripts upon the final acceptance of the paper. This timeline allows us to carefully clean and document the code for ease of use. Experimental settings, hyperparameters, and implementation details are described in Section˜4 and Section˜A.1 to facilitate verification prior to code release.
Appendix A Appendix
A.1 Urban scene reconstruction
A.1.1 NuPlan pose correction
We propose a pose correction method based on LiDAR point cloud registration in Section˜3.1.1. As shown in Figure˜7, point cloud registration effectively corrects pose calibration errors. This leads to improved scene reconstruction quality, which is reflected in the increased PSNR, as presented in Table˜4.
@@ A.1.1 NuPlan pose correction
A.1.2 Camera images reconstruction
@@
We use StreetGaussians (Yan et al., 2024b) to reconstruct the scene and perform additional processing on the nuPlan dataset. StreetGaussians (Yan et al., 2024b) uses manual pose annotation of dynamic objects to distinguish between static background and moving objects. Dynamic objects are reconstructed in their respective centered canonical spaces and then placed into the background scene space during the rendering process based on the known poses.
As shown in Figure˜8 (a), the camera images in nuPlan exhibit significant barrel distortion. Since 3DGS relies on image-based rendering and does not render rays corresponding to each pixel as NeRF does, such distortion adversely affects the reconstruction of scene geometry. To mitigate this issue, we applied a distortion correction to the images. Furthermore, due to varying exposure levels across the cameras in the nuPlan dataset, the same object exhibits substantial color differences when viewed from different cameras, as shown in Figure˜8 (b), leading to ambiguities in the color information of Gaussian primitives. To address this, we introduced a learnable affine transformation $\{{\bm{A}}_{i},{\bm{t}}_{i}\}$ for each camera $i$ to individually calibrate the image colors:
@@
where $i$ represents the camera index, ${\bm{C}}_{i}$ denotes the original rendered RGB map, and $\tilde{{\bm{C}}}_{i}$ refers to the RGB map after exposure transformation. These processing methods improve the quality of scene reconstruction, as shown in the ablation study in Table˜4.
@@
We compared our optimized reconstruction results with the state-of-the-art methods (Yan et al., 2024b; Chen et al., 2024; 2023) on 6 Navsim (Dauner et al., 2024) sequences, as shown in Figure˜3 and Table˜5, and observed a notable improvement in the reconstruction performance on nuPlan.
Table 5: Comparison with state-of-the-art camera images reconstruction methods in nuPlan.
[…]
StreetGaussians (Yan et al., 2024b) 26.02 0.798 0.179
Omnire (Chen et al., 2024) 26.89 0.837 0.165
PVG (Chen et al., 2023) 28.32 0.854 0.176
RealEngine(Ours) 29.67 0.897 0.093
@@
Although existing urban reconstruction methods perform excellently in synthesizing novel viewpoints for recorded trajectories, they face challenges when handling new trajectories in closed-loop simulations due to the limited viewpoints of driving videos and the vastness of the driving environment. To address this challenge, we employ DriveX (Yang et al., 2025) to enhance the reconstructed scene, which utilizes video generative prior (Yu et al., 2024a) to optimize the scene model across multiple trajectories. As shown in Figure˜9, our final scene model is capable of generating high-fidelity virtual driving environments beyond the recorded trajectories, enabling free-form trajectory driving simulations. We additionally conduct the evaluation protocol described in DriveX (Yang et al., 2025) to measure lateral offset on the nuPlan dataset. The results reported in Table˜6 also shows that our method is capable of synthesizing high-quality sensor data even when the ego-vehicle deviates from the original trajectory.
@@
Figure 9: Lane change reconstruction. We render the camera images by shifting the driving perspective 3 meters to the left and right. Due to the limited viewpoint of driving videos and the vastness of the driving environment, large translations from driving perspectives lead to a significant decline in reconstruction quality. To address this, we use DriveX (Yang et al., 2025), which leverages video generative prior to optimize the scene Gaussian primitives, resulting in improved reconstruction quality even with large viewpoint translations.
Table 6: Lane change reconstruction. We conduct the evaluation protocol described in DriveX (Yang et al., 2025) to measure lateral offset on the nuPlan dataset.
$\pm$0m (recorded) $\pm$1m $\pm$2m $\pm$3m
[…]
PVG (Chen et al., 2023) 28.32 0.854 62.71 95.34 130.49
StreetGaussian (Yan et al., 2024b) 26.02 0.798 59.48 87.74 103.30
DriveX (our use) (Yang et al., 2025) 29.67 0.897 52.59 77.27 86.84
A.1.3 LiDAR reconstruction
We use GS-LiDAR (Jiang et al., 2025) to reconstruct the LiDAR point clouds. GS-LiDAR (Jiang et al., 2025) introduces a novel panoramic rendering technique with explicit ray-splat intersection, guided by panoramic LiDAR supervision. This method employs 2D Gaussian primitives (Huang et al., 2024) with periodic vibration characteristics (Chen et al., 2023), allowing for precise geometric reconstruction of both static and dynamic elements in driving scenes.
The LiDAR point clouds in nuPlan (H. Caesar, 2021) is generated by merging point clouds from five LiDAR sensors, with non-uniform angular distributions across different scan lines. However, GS-LiDAR assumes that the LiDAR point cloud originate from a single laser scanner with uniformly spaced scan lines. To address this discrepancy, we reproject the merged point cloud onto the top LiDAR sensor to obtain the range map, selecting the point with the smaller depth when multiple points project to the same pixel. Additionally, we discard the non-uniform scan lines at both ends while retaining the central region, where the scan lines are evenly distributed.
As shown in Figure˜10, this processing method results in some loss of LiDAR information. However, since the driving agent (Chitta et al., 2023; Liao et al., 2025) converts the LiDAR point cloud into a 2-bin histogram over a 2D BEV grid with relatively low resolution, the missing information has a minimal impact on the driving agent. We evaluate the reconstruction performance on the same 6 sequences as the camera image reconstruction in Table˜7.
Figure 10: LiDAR reconstruction. Although reprojection may lead to some loss of LiDAR information, its impact on the low-resolution histogram used by the agent is minimal. Meanwhile, GS-LiDAR (Jiang et al., 2025) achieves high-quality reconstruction of the reprojected LiDAR data.